# Simple homotopy types of even dimensional manifolds

### v2

@@ 1.2. Simple homotopy manifold sets
    $|\mathcal{M}^{h}_{s}(M^{n}_{m})|\to\infty$ as $m\to\infty$, uniformly in $n$.
The ingredients of the proof will be discussed in Sections 1.4, 1.5, and 1.6. They rely on an analysis of the Whitehead group [Whitehead-she]. For a group $G$, Whitehead defined this group, $\operatorname{Wh}(G)$, and associated to a homotopy equivalence $f\colon X\rightarrow Y$ its Whitehead torsion $\tau(f)\in\operatorname{Wh}(\pi_{1}(Y))$. He proved the fundamental result that $\tau(f)=0$ if and only if $f$ is simple.
@@
First we will prove Theorem E (see Section 1.4), which characterises simple homotopy manifold sets in a general setting. It identifies $\mathcal{M}^{h}_{s}(M)$ with the orbit set of an action of $\operatorname{hAut}(M)$ on a subgroup of $\operatorname{Wh}(\pi_{1}(M))$, which depends on the surgery obstruction map in the surgery exact sequence of $M$. This is analogous to how the set of manifolds homotopy equivalent to a given Poincaré complex can be computed as a quotient of its structure set. However, we prove (see Corollary 1.3) that for some fundamental groups, including $C_{\infty}\times C_{m}$, the relevant subgroup of $\operatorname{Wh}(\pi_{1}(M))$ depends only on $\pi_{1}(M)$ and the orientation character of $M$. To study this subgroup for $S^{1}\times L$, we will use representation theory and number theory, as explained in Section 1.6.
The action of a homotopy automorphism $g\in\operatorname{hAut}(M)$ on $\operatorname{Wh}(\pi_{1}(M))$ is determined by its Whitehead torsion $\tau(g)\in\operatorname{Wh}(\pi_{1}(M))$ and the induced homomorphism $\pi_{1}(g)\in\operatorname{Aut}(\pi_{1}(M))$ (see Definition 4.3). To understand this action for the manifolds $S^{1}\times L$, the key result is the following (see Theorem 6.6):
Theorem 1.2.
Every homotopy automorphism $f\colon S^{1}\times L\to S^{1}\times L$ is simple.
@@
This result explains why we are able to so accurately compute simple homotopy manifold sets for this class of manifolds. Note though that further work is still required, since the action of $\operatorname{hAut}(S^{1}\times L)$ on $\pi_{1}(S^{1}\times L)$ is nontrivial. See Section 1.5 for further details.
1.3. Comparison with $h$-cobordism
Recall that a cobordism $(W;M,N)$ of closed manifolds is an $h$-cobordism if the inclusion maps $M\to W$ and $N\to W$ are homotopy equivalences. The notion of an $h$-cobordism is central to manifold topology. Smale’s $h$-cobordism theorem [smale-pe, Smale-h-cob, Milnor-h-cob], together with its extensions to other categories and dimension 4 in [Stallings-Tata, Kirby-Siebenmann:1977-1, FQ], states that under 1.1, every simply-connected $h$-cobordism is $\operatorname{CAT}$-equivalent to the product $M\times I$. To generalise the $h$-cobordism theorem to non-simply connected manifolds, one also considers $s$-cobordisms. An $s$-cobordism $(W;M,N)$ is a cobordism for which the inclusions $M\to W$ and $N\to W$ are simple homotopy equivalences. The $\operatorname{CAT}$ $s$-cobordism theorem [Barden, Mazur:1963-1, Stallings-Tata, Kirby-Siebenmann:1977-1, FQ] states that under 1.1, $\operatorname{CAT}$ equivalence classes of $h$-cobordisms based on $M$ are in bijection with $\operatorname{Wh}(\pi_{1}(M))$, with the bijection given by taking the Whitehead torsion of the inclusion $M\to W$ (Theorem 3.1). In particular every $s$-cobordism is a product. This theorem underpins manifold classification in dimension at least $4$.
The s-cobordism theorem is one of the tools that we can use to construct homotopy equivalent but not simple homotopy equivalent manifolds (see Sections 3.1 and 4.2), and we note that the examples constructed this way are always h-cobordant. This leads us to consider two refined versions of the problem of finding homotopy but not simple homotopy equivalent manifolds, where the manifolds are also required to be h-cobordant, or required not to be h-cobordant. We introduce the corresponding variations of $\mathcal{M}^{h}_{s}(M)$:
$\mathcal{M}^{\operatorname{hCob}}_{s}(M)$$\coloneqq\left\{\operatorname{CAT}\text{ $n$-manifolds }N\mid N\text{ is $h$-cobordant to }M\right\}/\simeq_{s}$
$\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$$\coloneqq\left\{\operatorname{CAT}\text{ $n$-manifolds }N\mid N\simeq M\right\}/\langle\simeq_{s},\operatorname{hCob}\rangle$
where $\langle\simeq_{s},\operatorname{hCob}\rangle$ denotes the equivalence relation generated by simple homotopy equivalence and $h$-cobordism. We also note that the sets $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$ and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$ will arise naturally in the computation of $\mathcal{M}^{h}_{s}(M)$ in Theorem E below.
@@ 1.3. Comparison with $h$-cobordism
The next theorem shows that we obtain many examples where $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$ is nontrivial. Thus, even if we restrict to $h$-cobordant manifolds, many of the simple homotopy sets we consider remain large. As with $\mathcal{M}^{h}_{s}(M)$, the sets $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$ can be trivial, finite, or infinite, all for manifolds of the form $M=S^{1}\times L$, where $L$ is a lens space. Again using our number theoretic analysis of the Whitehead group, we determine precisely when each eventuality occurs, and we analyse the asymptotic behaviour.
Theorem C.
@@
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Let $M^{n}_{m}=S^{1}\times L$, where $L$ is an $(n-1)$-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, $m\geq 2$. Then
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|$ only depends on $n$ and $m$, but it is independent of the choice of $L$ or $\operatorname{CAT}$;
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$;
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free;
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{s}(M^{n}_{m})|\to\infty$ as $m\to\infty$ uniformly in $n$.
On the other hand, we also discover a wealth of interesting behaviour when we also factor out by $h$-cobordism.
Theorem D.
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Let $M^{n}_{m}=S^{1}\times L$, where $L$ is an $(n-1)$-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, $m\geq 2$. Then
@@
    $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|$ only depends on $n$ and $m$, but it is independent of the choice of $L$ or $\operatorname{CAT}$;
@@
    $\liminf_{m\to\infty}\Bigl{(}\sup_{n}|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|\Bigr{)}=1$;
@@
    $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|<\infty$ for all $n$ and $m$;
@@
    $\limsup_{m\to\infty}\Bigl{(}\inf_{n}|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|\Bigr{)}=\infty$.
@@
Note that part (b) says that there are infinitely many $m$ with $|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|=1$ for every $n$, while part (d) says that $\inf_{n}|\mathcal{M}^{h}_{s,\operatorname{hCob}}(M^{n}_{m})|$ is unbounded in $m$.
1.4. Characterisation of simple homotopy manifold sets
In Theorem E below, we characterise the simple homotopy manifold sets $\mathcal{M}^{h}_{s}(M)$, $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$, and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$, for a $\operatorname{CAT}$ $n$-manifold $M$, in terms of the Whitehead group, the homotopy automorphisms of $M$, and the surgery obstruction map. All our main theorems above arise by applying these characterisations to the manifolds $S^{1}\times L$ and computing the objects that appear in Theorem E.
An orientation character $w\colon G\to C_{2}$ determines an involution $x\mapsto\overline{x}$ on the Whitehead group $\operatorname{Wh}(G)$ (see Section 2.2), and we will write $\operatorname{Wh}(G,w)$ to specify that $\operatorname{Wh}(G)$ is equipped with this involution. Define:
$\mathcal{J}_{n}(G,w)$ =$=\left\{y\in\operatorname{Wh}(G,w)\mid y=-(-1)^{n}\overline{y}\right\}\leq\operatorname{Wh}(G,w),$
$\mathcal{I}_{n}(G,w)$ =$=\left\{x-(-1)^{n}\overline{x}\mid x\in\operatorname{Wh}(G,w)\right\}\leq\mathcal{J}_{n}(G,w).$
The Tate cohomology group $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is canonically identified with $\mathcal{J}_{n}(G,w)/\mathcal{I}_{n}(G,w)$ (see Proposition 9.5), and we will denote the quotient map by
If $w\equiv 1$ is the trivial character, then we omit it from the notation.
There is an action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$ (as a set) such that if $f\colon N\rightarrow M$ is a homotopy equivalence and $g\in\operatorname{hAut}(M)$, then $\tau(f)^{g}=\tau(g\circ f)$. Let $q\colon\operatorname{Wh}(G,w)\rightarrow\operatorname{Wh}(G,w)/\operatorname{hAut}(M)$ denote the quotient map, let $\varrho\colon\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))\rightarrow L_{n}^{s}(\mathbb{Z}G,w)$ be the homomorphism from the Ranicki-Rothenberg exact sequence (3.1) (see [Shaneson-GxZ], [Ranicki-80-I, §9]), and let $\sigma_{s}\colon\mathcal{N}(M)\rightarrow L_{n}^{s}(\mathbb{Z}G,w)$ be the surgery obstruction map (see Section 3.2). The following theorem is the basis of our main results (see Theorems 4.11 and 4.16).
Theorem E.
Let $M$ be a $\operatorname{CAT}$ $n$-manifold with fundamental group $G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, satisfying 1.1. There is a commutative diagram
[…]
The subset $\mathcal{J}_{n}(G,w)$ is invariant under the action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$, and there is an induced action on $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$. In the statement above, the subsets
$(\varrho\circ\pi)^{-1}(\operatorname{Im}\sigma_{s})\subseteq\mathcal{J}_{n}(G,w)\quad\text{ and }\quad\varrho^{-1}(\operatorname{Im}\sigma_{s})\subseteq\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$
are both invariant under the induced actions of $\operatorname{hAut}(M)$. Note that if the subset $\mathcal{I}(G,w)$ is invariant under the action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$, then $q(\mathcal{I}_{n}(G,w))=\mathcal{I}_{n}(G,w)/\operatorname{hAut}(M)$. It follows from the exact sequence (3.1) that the image of the homomorphism $\psi\colon L_{n+1}^{h}(\mathbb{Z}G,w)\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is contained in $\varrho^{-1}(\operatorname{Im}\sigma_{s})$ and so we obtain the following corollary.
Corollary 1.3.
Let $M$ be a $\operatorname{CAT}$ $n$-manifold with fundamental group $G$ and orientation character $w\colon G\rightarrow\left\{\pm 1\right\}$, satisfying 1.1. If $\psi\colon L_{n+1}^{h}(\mathbb{Z}G,w)\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ is surjective, then we have
$\mathcal{M}^{h}_{s}(M)\cong\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M),\quad\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)\cong\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M).$
The advantage of Corollary 1.3 is that $\mathcal{J}_{n}(G,w)$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))$ (and, more generally, $\operatorname{Im}\psi$) only depend on $G$ and $w$, while $\operatorname{Im}\sigma_{s}$ depends on $M$ a priori. This will allow us to apply Corollary 1.3 by separately analysing the involution on $\operatorname{Wh}(G,w)$ and the homotopy automorphisms of $M$. Note also that, in cases where $M$ can be regarded as a manifold in multiple categories, the sets $q(\mathcal{I}_{n}(G,w))$, $\mathcal{J}_{n}(G,w)/\operatorname{hAut}(M)$ and $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(G,w))/\operatorname{hAut}(M)$ do not depend on the choice of $\operatorname{CAT}$.
To prove Theorem E, first we observe that if $N$ is a manifold homotopy equivalent to $M$, then the set of Whitehead torsions of all possible homotopy equivalences $f\colon N\to M$ forms an orbit of the action of $\operatorname{hAut}(M)$ on $\operatorname{Wh}(G,w)$. By considering the different restrictions and equivalence relations on these manifolds $N$, we obtain maps from the various simple homotopy manifold sets of $M$ to subsets/subquotients of $\operatorname{Wh}(G,w)/\operatorname{hAut}(M)$. We then verify that these maps are injective, and the remaining task is to determine their images.
For $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$, we use the $s$-cobordism theorem. For every $x\in\operatorname{Wh}(G,w)$ there is an $h$-cobordism $(W;M,N)$ with Whitehead torsion $x$, and then $\tau(f)=-x+(-1)^{n}\overline{x}$ for the induced homotopy equivalence $f\colon N\rightarrow M$ (see Proposition 2.38). This shows that every element of $\mathcal{I}_{n}(G,w)$ can be realised as the torsion of a homotopy equivalence $N\rightarrow M$ for some $N$ that is $h$-cobordant to $M$. For $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$, we use the surgery exact sequence combined with the exact sequence (3.1) to describe the set of values of $\pi(\tau(f))$ for all homotopy equivalences $f\colon N\to M$ (see Proposition 3.6). Finally, the characterisation of $\mathcal{M}^{h}_{s}(M)$ is obtained by combining the results on $\mathcal{M}^{\operatorname{hCob}}_{s}(M)$ and $\mathcal{M}^{h}_{s,\operatorname{hCob}}(M)$.
1.5. Outline of the proof of Theorems B, C and D
Recall that a lens space $L\coloneqq S^{2k-1}/C_{m}$ of dimension $2k-1\geq 3$ is a quotient of $S^{2k-1}$ by a free action of a finite cyclic group $C_{m}$ for some $m\geq 2$. The action is determined by a $k$-tuple of integers $(q_{1},\dots,q_{k})$ with $q_{j}$ coprime to $m$ for all $j$. The fundamental group is $\pi_{1}(L)\cong C_{m}$.
@@ 1.5. Outline of the proof of Theorems B, C and D
First we consider $\mathcal{M}^{\operatorname{hCob}}_{s}(S^{1}\times L)$ and the proof of Theorem B (and hence Theorem A). The proof is based on applying Corollary 1.3 to $S^{1}\times L$. This is possible, because the map $\psi\colon L_{n+1}^{h}(\mathbb{Z}[C_{\infty}\times C_{m}])\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$ is surjective for $n=2k$ (see Proposition 3.12), and we get that $\mathcal{M}^{h}_{s}(S^{1}\times L)\cong\mathcal{J}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(S^{1}\times L)$. For the homotopy automorphisms of $S^{1}\times L$, we will use the following, expanding on Theorem 1.2 (see Theorems 6.6 and 6.5).
@@
    Every homotopy automorphism $f\colon S^{1}\times L\to S^{1}\times L$ is simple.
@@
    If $\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(C_{\infty}\times C_{m})$ is the map given by taking the induced automorphism on the fundamental group, then $\operatorname{Im}(\pi_{1})=\left\{\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)\in\operatorname{Aut}(C_{\infty}\times C_{m})\mid c^{k}\equiv\pm 1\text{ mod }m\right\}$.
Part (a) implies that the action of $\operatorname{hAut}(S^{1}\times L)$ on $\operatorname{Wh}(C_{\infty}\times C_{m})$ factors through the action of $\operatorname{Aut}(C_{\infty}\times C_{m})$, which acts on $\operatorname{Wh}(C_{\infty}\times C_{m})$ via automorphisms, using the functoriality of $\operatorname{Wh}$ (see Definitions 4.3 and 4.9). Therefore every orbit has cardinality at most $|\operatorname{Aut}(C_{\infty}\times C_{m})|<2m^{2}$, and $|\mathcal{J}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(S^{1}\times L)|=1$ if and only if $\mathcal{J}_{n}(C_{\infty}\times C_{m})=0$. Moreover, since it follows from part (b) that $\operatorname{Im}(\pi_{1}\colon\operatorname{hAut}(S^{1}\times L)\rightarrow\operatorname{Aut}(C_{\infty}\times C_{m}))$ is independent of the choice of the integers $(q_{1},\dots,q_{k})$, the same is true for $\mathcal{J}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(S^{1}\times L)$, proving part (a) of Theorem B.
So it remains to study the involution on $\operatorname{Wh}(C_{\infty}\times C_{m})$ and prove the corresponding statements about $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|$. First, the fundamental theorem for $K_{1}(\mathbb{Z}C_{m}[t,t^{-1}])$ gives rise to a direct sum decomposition (see Theorem 5.9):
$\operatorname{Wh}(C_{m}\times C_{\infty})\cong\operatorname{Wh}(C_{m})\oplus\widetilde{K}_{0}(\mathbb{Z}C_{m})\oplus NK_{1}(\mathbb{Z}C_{m})^{2}$
@@
$\mathcal{J}_{n}(C_{\infty}\times C_{m})\cong\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid x=-\overline{x}\}\oplus NK_{1}(\mathbb{Z}C_{m}).$
In this decomposition the first component is always finite (see Lemma 5.12), so Theorem B (c) is proved by the following (see Theorems 5.8 and 5.7).
@@
For the remaining parts (b) and (d) of Theorem B, we need to study the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ and prove the analogous statements about $\{x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid x=-\overline{x}\}$ in the case where $m$ is square-free. This will be achieved using methods from algebraic number theory, and discussed further in Section 1.6 (see Theorem 1.6 below).
Now we consider $\mathcal{M}^{\operatorname{hCob}}_{s}(S^{1}\times L)$ and the proof of Theorem C. According to Theorem E we need to analyse $\mathcal{I}_{n}(C_{\infty}\times C_{m})$. We can do this similarly to $\mathcal{J}_{n}(C_{\infty}\times C_{m})$. The former has the following decomposition (see Proposition 5.10):
$\mathcal{I}_{n}(C_{\infty}\times C_{m})\cong\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}\oplus NK_{1}(\mathbb{Z}C_{m}).$
It follows from Theorem 1.4 that $\mathcal{I}_{n}(C_{\infty}\times C_{m})$ is invariant under the action of $\operatorname{hAut}(S^{1}\times L)$, so $q(\mathcal{I}_{n}(C_{\infty}\times C_{m}))$ can be expressed as $\mathcal{I}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(S^{1}\times L)$. Thus by Theorem E we have $\mathcal{M}^{\operatorname{hCob}}_{s}(S^{1}\times L)\cong\mathcal{I}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(S^{1}\times L)$. By Theorems 1.4 and 1.5, together with Theorem 1.7 below, this leads to Theorem C.
Finally, we consider $\mathcal{M}^{h}_{s,\operatorname{hCob}}(S^{1}\times L)$ and the proof of Theorem D. By Corollary 1.3, this is isomorphic to $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))/\operatorname{hAut}(S^{1}\times L)$. The decomposition of $\operatorname{Wh}(C_{\infty}\times C_{m})$ induces an isomorphism $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))\cong\widehat{H}^{n+1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))$. Hence, as above, it remains to consider the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$, but the behaviour of $\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$ is different from $\mathcal{J}_{n}(C_{\infty}\times C_{m})$ and $\mathcal{I}_{n}(C_{\infty}\times C_{m})$, in particular it is always finite. The key algebraic input comes from Theorem 1.8 below.

### v3

@@ 1.3. Computations of simple homotopy manifold sets
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|\to\infty$ as $m\to\infty$, uniformly in $n$.
@@
The proof is an application of Theorem B. This theorem identifies $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M)$ with the orbit set of an action of $\operatorname{hAut}(M)$ on a subgroup of $\operatorname{Wh}(\pi_{1}(M))$, which depends on the surgery obstruction map in the surgery exact sequence of $M$. However as in Corollary 1.3, for some fundamental groups, including $C_{\infty}\times C_{m}$, the relevant subgroup of $\operatorname{Wh}(\pi_{1}(M))$ depends only on $\pi_{1}(M)$ and the orientation character of $M$. To study this subgroup for $S^{1}\times L$, we use representation theory and number theory, as explained in Section 1.5.
@@
The action of an element $g\in\operatorname{hAut}(M)$ on $\operatorname{Wh}(\pi_{1}(M))$ is determined by its Whitehead torsion $\tau(g)\in\operatorname{Wh}(\pi_{1}(M))$ and the induced homomorphism $\pi_{1}(g)\in\operatorname{Aut}(\pi_{1}(M))$ (see Definition 4.3). It turns out that computing this action is particularly tractable for $S^{1}\times L$; see Theorem 1.5 for more details. This explains why we are able to so accurately compute simple homotopy manifold sets for this class of manifolds.
@@
The next theorem gives examples where $\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M)$ is nontrivial. Thus, even if we restrict to $h$-cobordant manifolds, many of the simple homotopy manifold sets we consider remain large.
Theorem D.
@@
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Let $L$ be an (n1)(n-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, for some $m\geq 2$, and let $M^{n}_{m}=S^{1}\times L$. Then
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|$ only depends on $n$ and $m$, but it is independent of the choice of $L$ or $\operatorname{CAT}$;
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|=1$ if and only if $m\in\{2,3,5,6,7,10,11,13,14,15,17,19,29\}$;
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ if and only if $m$ is not square-free;
@@
    $|\mathcal{M}^{\operatorname{hCob}}_{\mathrm{s}}(M^{n}_{m})|\to\infty$ as $m\to\infty$ uniformly in $n$.
Finally, we discover a wealth of interesting behaviour when we also factor out by $h$-cobordism.
Theorem E.
Let $n\geq 4$ be even and let $\operatorname{CAT}$ be as in 1.1. Let $L$ be an (n1)(n-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, for some $m\geq 2$, and let $M^{n}_{m}=S^{1}\times L$. Then
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|$ only depends on $n$ and $m$, but it is independent of the choice of $L$ or $\operatorname{CAT}$;
@@
    $\liminf_{m\to\infty}\Bigl(\sup_{n}|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|\Bigr)=1$;
@@
    $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|<\infty$ for all $n$ and $m$;
@@
    $\limsup_{m\to\infty}\Bigl(\inf_{n}|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|\Bigr)=\infty$.
@@
Part (b) says that there are infinitely many $m$ with $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|=1$ for every $n$, while part (d) says that $\inf_{n}|\mathcal{M}^{\mathrm{h}}_{\mathrm{s},\operatorname{hCob}}(M^{n}_{m})|$ is unbounded in $m$.
1.4. Outline of the proof of Theorem A
@@
Let $n=2k\geq 4$ be even and let $M^{n}_{m}=S^{1}\times L$, where $L$ is an (n1)(n-1)-dimensional lens space with $\pi_{1}(L)\cong C_{m}$, $m\geq 2$. It suffices to show that $|\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})|=\infty$ for some $m\geq 2$. The proof is based on applying Corollary 1.3 to $M^{n}_{m}$. We show that $\psi\colon L_{n+1}^{\mathrm{h}}(\mathbb{Z}[C_{\infty}\times C_{m}])\rightarrow\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))$ is surjective for $n$ even in Proposition 7.1, and so Corollary 1.3 implies that $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})\cong\mathcal{J}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(M^{n}_{m})$. For the homotopy automorphisms of $M^{n}_{m}$, we use the following (see Theorems 6.2 and 6.6).
@@
    Every homotopy automorphism $f\colon M^{n}_{m}\to M^{n}_{m}$ is simple.
@@
    If $\pi_{1}\colon\operatorname{hAut}(M^{n}_{m})\rightarrow\operatorname{Aut}(C_{\infty}\times C_{m})$ is the map given by taking the induced automorphism on the fundamental group, then $\operatorname{Im}(\pi_{1})=\left\{\left(\begin{smallmatrix}a&b\\ 0&c\end{smallmatrix}\right)\in\operatorname{Aut}(C_{\infty}\times C_{m})\mid c^{k}\equiv\pm 1\text{ mod }m\right\}$.
Part (a) implies that the action of $\operatorname{hAut}(M^{n}_{m})$ on $\operatorname{Wh}(C_{\infty}\times C_{m})$ factors through the action of $\operatorname{Aut}(C_{\infty}\times C_{m})$, which acts on $\operatorname{Wh}(C_{\infty}\times C_{m})$ via automorphisms, using the functoriality of $\operatorname{Wh}$ (see Definitions 4.3 and 4.9). Therefore every orbit has cardinality bounded above by $|\hskip-1.42262pt\operatorname{Aut}(C_{\infty}\times C_{m})|<2m^{2}$, and $|\mathcal{J}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(M^{n}_{m})|=1$ if and only if $\mathcal{J}_{n}(C_{\infty}\times C_{m})=0$.
So to prove Theorem A, and more generally Theorem C, it remains to study the involution on $\operatorname{Wh}(C_{\infty}\times C_{m})$ and prove the corresponding statements about $|\mathcal{J}_{n}(C_{\infty}\times C_{m})|$. First, the fundamental theorem for $K_{1}(\mathbb{Z}C_{m}[t,t^{-1}])$ gives rise to a direct sum decomposition (see Theorem 5.9):
$\operatorname{Wh}(C_{\infty}\times C_{m})\cong\operatorname{Wh}(C_{m})\oplus\widetilde{K}_{0}(\mathbb{Z}C_{m})\oplus NK_{1}(\mathbb{Z}C_{m})^{2}$
@@
$\mathcal{J}_{n}(C_{\infty}\times C_{m})\cong\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid y=-\overline{y}\}\oplus NK_{1}(\mathbb{Z}C_{m}).$
@@
In particular $\mathcal{J}_{n}(C_{\infty}\times C_{m})$ is also infinite in the latter case. Since each orbit of $\operatorname{hAut}(M^{n}_{m})$ is finite, there are infinitely many orbits. Since $\mathcal{M}^{\mathrm{h}}_{\mathrm{s}}(M^{n}_{m})\cong\mathcal{J}_{n}(C_{\infty}\times C_{m})/\operatorname{hAut}(M^{n}_{m})$, we obtain Theorem A by taking $m$ not square-free, e.g. $m=4$.
The proofs of Theorems D and E are similarly approached using Corollary 1.3 and Theorem 1.5, as well as the following decompositions (see Propositions 5.10 and 5.11):
$\mathcal{I}_{n}(C_{\infty}\times C_{m})\cong\{x-\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}\oplus NK_{1}(\mathbb{Z}C_{m})$
$\widehat{H}^{n+1}(C_{2};\operatorname{Wh}(C_{\infty}\times C_{m}))\cong\widehat{H}^{n+1}(C_{2};\widetilde{K}_{0}(\mathbb{Z}C_{m}))\cong\frac{\{y\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\mid\overline{y}=-(-1)^{n}y\}}{\{x-(-1)^{n}\overline{x}\mid x\in\widetilde{K}_{0}(\mathbb{Z}C_{m})\}}.$
The finiteness results follow from Theorem 1.6 combined with the fact that $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ is finite (see Proposition 8.4). However, the results on triviality and asymptotic behaviour of the manifold sets requires a deeper analysis of the involution on $\widetilde{K}_{0}(\mathbb{Z}C_{m})$ (see Theorems 1.7 and 1.8 below).